# Pitch: AI-assisted Tagging of Deepfake Audio Calls using Challenge-Response

### v3

@@ 6. Discussion
Practical Design Recommendations. To ensure a practical implementation of the ‘Deepfake-Likely’ tagging system, we offer the following guidelines:
On the caller’s side:
Utilize single, succinct sentences for tasks.
Enhance compliance with high-quality instructional samples.
For playback, employ an automated sound for ease of use.
On the receiver’s side:
Providing a rationale to back prediction has nominal benefits.
Limitations and Future Work. Our study has several limitations that provide directions for future research:
Comprehensive challenge evaluation: We did not assess all sub-categories in our challenge taxonomy, particularly those that might prove more relevant in a multilingual context, eg, roll R.
Testing against COTS deepfake generators: Pitch has not been tested against commercial off-the-shelf (COTS) deepfake generators like ElevenLabs.
Advanced machine learning models: Future work could utilize more sophisticated and recent foundation models to enhance cross-lingual machine detection accuracy.
Dataset diversity: Expanding our dataset to capture greater linguistic diversity, including speech variations like lisps, could improve the system’s robustness.
Inclusivity: Developing a more inclusive system that can calibrate detection thresholds for individuals with known speech impediments is an important addition.
@@ 7. Conclusion
This study presents a novel challenge-response approach to enhance both human and machine detection of interactive deepfake audio calls. We developed a comprehensive taxonomy of audio challenges, identifying 10 robust instances from 20 candidates, and provide a diverse open-source dataset. These challenges in the dataset effectively degraded deepfake quality, thereby improving machine detection capabilities and enabling humans to discern deepfakes, even against a challenging dataset baseline.
By fostering a synergistic relationship between humans and machines, we achieved a balance that capitalizes on human oversight and machine efficiency. This complementary approach has laid the foundation for an innovative tagging system aimed at pre-screening calls, marking a significant advancement in audio deepfake detection that operates independently of target identity.
@@
| General Sentences |
| The quick brown fox jumps over the lazy dog. |
| The swift cheetah is an elegant and extraordinary creature. |
| The rugged path wound around the steep hill to the quaint village. |
| Gazing at the azure sky in July, the dreamy poet wrote verses. |
| The glum monk thought deeply about life’s complex queries. |
| Questions |
| Does the quick brown fox jump over the lazy dog? |
| Is the swift cheetah an elegant and extraordinary creature? |
| Does the rugged path wind around the steep hill to the quaint village? |
| Did the glum monk think deeply about life’s complex queries? |
| Foreign Sentences |
| Finnish: ’En ymmarra mitaan mita tarkoitat’ |
| French: ’Le coeur a ses raisons que la raison ne connait point.’ |
| Spanish: ’El rapido zorro marron salta sobre el perro perezoso.’ |
| Chinese: ’Feng yu zhi hou, bi jian cai hong.’ |
| Russian: ’Na dvore trava, na trave drova.’ |
| Arabic: ’Al-hayatu jamilatun ala al-raghmi min kulli shay’. |
| Hindi: ’Bandar bhagate hue chhat par chadh gaya.’ |
| Japanese: ’Sakura no hana ga haru no kaze ni mau.’ |
| Swahili: ’Simba anapenda kupumzika chini ya mti mkubwa.’ |
A.1. Data Collection Instructions
@@
| Challenge | DF1 | DF2 | VC1 | VC2 | ||

### v4

@@ 6. Discussion
Accessibility and Inclusivity Considerations. Pitch’s vocal challenges could present barriers for speech-impaired individuals as they might be incorrectly flagged as suspicious. We propose an accommodation pathway where users can self-identify speech impairments, triggering automatic human review with alternative authentication methods (pre-registered devices/emails). To prevent exploitation of this pathway, machine evaluation should still run in parallel, providing reviewers with supplementary data for informed decision-making.
Also, though our evaluation focused on English speakers, Pitch’s taxonomy is fundamentally language-agnostic and extends readily to multilingual implementations. Notably, accent diversity among non-native speakers strengthens the system’s effectiveness, as acoustic variations in accented speech create additional complexity for voice conversion systems trained on homogeneous data.
Hence, Pitch balances accessibility needs with security requirements, while being inclusive and maintaining verification integrity.
Limitations and Future Work.
Comprehensive challenge evaluation: Our study did not assess all sub-categories in our challenge taxonomy, particularly those that might prove more relevant in a multi-lingual context.
Testing against COTS deepfake generators: Pitch has not been tested against commercial off-the-shelf deepfake generators.
Advanced ML models: Utilize more sophisticated and recent foundation models to enhance cross-lingual machine evaluation.
@@ 7. Conclusion
This study presents a novel challenge-response approach, Pitch, to detect interactive deepfake audio calls in real-world authentication scenarios. We developed a comprehensive taxonomy of audio challenges, identifying 10 robust instances from 20 candidates, and created a diverse open-source dataset of 18,600 original and 1.6 million deepfake samples. These challenges effectively degraded deepfake quality, improving machine detection capabilities from 56.0% to 88.7% AUROC and enabling humans to discern deepfakes with 72.6% accuracy, even against a challenging dataset baseline.
By fostering a synergistic relationship between humans and machines, we achieved a balanced framework that capitalizes on human oversight and machine efficiency. Our collaborative approach reached 84.5% detection accuracy while maintaining human decision control in 56% of cases—a significant improvement over individual performance. This methodology provides a practical solution that requires no prior enrollment or reference samples, laying the foundation for robust defensive measures against increasingly sophisticated voice-based social engineering attacks across government, healthcare and finance sectors.
@@
Glossary:
A1.
Human Cognitive Bias and Susceptibility against Deepfakes.
A2.
Instructions provided to participants for data collection.
A3.
Deployment Considerations.
A4.
Training of Automated Challenge Compliance Detectors.
A5.
Instruments for Human Evaluation of Deepfakes.
A.1. Cognitive Bias and Human Susceptibility
A.2. Data Collection Instructions
@@
| General Sentences |
| The quick brown fox jumps over the lazy dog. |
| The swift cheetah is an elegant and extraordinary creature. |
| The rugged path wound around the steep hill to the quaint village. |
| Gazing at the azure sky in July, the dreamy poet wrote verses. |
| The glum monk thought deeply about life’s complex queries. |
| Questions |
| Does the quick brown fox jump over the lazy dog? |
| Is the swift cheetah an elegant and extraordinary creature? |
| Does the rugged path wind around the steep hill to the quaint village? |
| Did the glum monk think deeply about life’s complex queries? |
| Foreign Sentences |
| Finnish: ’En ymmarra mitaan mita tarkoitat’ |
| French: ’Le coeur a ses raisons que la raison ne connait point.’ |
| Spanish: ’El rapido zorro marron salta sobre el perro perezoso.’ |
| Chinese: ’Feng yu zhi hou, bi jian cai hong.’ |
| Russian: ’Na dvore trava, na trave drova.’ |
| Arabic: ’Al-hayatu jamilatun ala al-raghmi min kulli shay’. |
| Hindi: ’Bandar bhagate hue chhat par chadh gaya.’ |
| Japanese: ’Sakura no hana ga haru no kaze ni mau.’ |
| Swahili: ’Simba anapenda kupumzika chini ya mti mkubwa.’ |
A.3. Practical Design Recommendations
To ensure a practical implementation of the ‘Deepfake-Likely’ tagging system, we offer the following guidelines:
On the caller’s side:
Utilize single, succinct sentences for tasks.
Enhance compliance with high-quality instructional samples.
For playback, employ an automated sound for ease of use.
On the receiver’s side:
Providing a rationale to back prediction has nominal benefits.
A.4. Challenge Compliance Detection Architecture
Our challenge compliance detection system employs a series of binary classifiers based on the Wav2Vec2 architecture (Baevski et al., 2020).
Wav2Vec2 Model Architecture. Wav2Vec2 is a self-supervised framework for speech representation learning that consists of a multi-layer convolutional feature encoder followed by a transformer context network. The architecture processes raw audio waveforms directly, eliminating the need for feature engineering.
For our implementation, we used the public wav2vec2-base model pretrained on 960 hours of LibriSpeech data, which was further pretrained on 50,000 hours of unlabeled speech from Libri-Light. This extensive pretraining enables the model to learn robust speech representations across diverse acoustic conditions.
Fine-tuning for Challenge Compliance Detection. We adapted the pretrained Wav2Vec2 model for challenge compliance detection through the following modifications:
Classification Head: We replaced the original masked language modeling head with a binary classification layer consisting of a linear projection (768 → 256), followed by layer norm, ReLU activation, dropout (p=0.1), and a final linear layer (256 → 2).
Input Processing: Audio samples were resampled to 16kHz and normalized to zero mean and unit variance. For each 4.5±1.1s challenge sample, we extracted fixed-length segments of 4 seconds from the center of the utterance, with zero-padding for shorter segments.
We implemented separate binary classifiers for each of the 18 challenges, with each classifier distinguishing between its specific challenge and regular speech (challenge #0). The remaining 2 challenges (foreign words and questions) were evaluated using heuristic approaches based on speech transcription as described in Table A.2.
| Baseline | Val Acc. | Test Acc. | AUC | Precision | Recall | F1 | ||
| #1 | Static Mouth | 85.9 | 88.2 | 96.6 | 0.89 | 0.87 | 0.88 | |
| #2 | Cup mouth | 86.9 | 89.5 | 97.1 | 0.95 | 0.84 | 0.89 | |
| #3 | Whisper | 94.9 | 90.9 | 93.1 | 0.92 | 0.9 | 0.91 | |
| #4 | Hold nose | 43.4 | 90.0 | 92.8 | 0.94 | 0.85 | 0.90 | |
| #5 | High Pitch | 86.9 | 81.8 | 91.2 | 0.95 | 0.67 | 0.79 | |
| #6 | Low Pitch | 73.7 | 77.3 | 85.9 | 0.75 | 0.83 | 0.78 | |
| #7 | Sing | 87.9 | 90.0 | 96.8 | 0.89 | 0.91 | 0.9 | |
| #8 | Speak Loudly | 86.9 | 70.5 | 74.8 | 0.73 | 0.65 | 0.69 | |
| #9 | Speak softly | 85.9 | 79.1 | 84.7 | 0.76 | 0.85 | 0.8 | |
| #10 | Read quickly | 98.0 | 83.2 | 90.7 | 0.92 | 0.73 | 0.81 | |
| #11 | Read Slowly | 89.9 | 79.1 | 87.5 | 0.8 | 0.77 | 0.79 | |
| #13 | Accent | 72.7 | 79.5 | 85.9 | 0.81 | 0.77 | 0.79 | |
| #14 | Emotions | 89.9 | 88.6 | 95.6 | 0.98 | 0.79 | 0.87 | |
| #16 | Cough/Whistle | 90.9 | 96.7 | 99.3 | 0.82 | 0.82 | 0.82 | |
| #17 | Clap | 84.7 | 94.1 | 98.8 | 0.95 | 0.93 | 0.94 | |
| #18 | Cross-talk | 91.5 | 98.8 | 100.0 | 1.0 | 0.96 | 0.98 | |
| #19 | Instr. Playback | 95.7 | 98.1 | 99.6 | 0.98 | 0.96 | 0.97 | |
| #20 | Lyric Playback | 100.0 | 98.1 | 100.0 | 0.94 | 1.0 | 0.97 | |
| Average (across all) | 86.0 | 87.5 | 93.1 | 0.89 | 0.82 | 0.85 | ||
| Heuristic-based Compliance Check | ||||||||
| #15 | Question | 92.9 | Transcribed sentence has a question mark | |||||
Training Protocol. The training protocol for our challenge compliance classifiers consisted of:
Dataset Partitioning: We partitioned the dataset by subject identity for 5-fold cross-validation, using 90 identities for training/validation and reserving 10 identities for a final hold-out test set. This identity-based partitioning ensures no speaker overlap between training and evaluation sets, testing the model’s generalization to unseen speakers.
Optimization: Models were fine-tuned using AdamW optimizer with a learning rate of 5e-5, , , weight decay of 0.01, and a linear warmup for the first 10% of training steps followed by linear decay.
Training Configuration: Each model was trained for 30 epochs with a batch size of 64.
Regularization: We applied dropout (p=0.1) to the classification head and the attention layers, along with layer drop (p=0.05) for transformer layers to prevent overfitting.
Model Selection: After each epoch, we monitored validation accuracy and retained the best-performing model checkpoint based on this metric.
Implementation Performance. Our Wav2Vec2-based challenge compliance detection system achieved:
Mean validation accuracy of 87.5% across all 18 chal. classifiers
Mean test accuracy of 83.6% ± 8.0% on the hold-out test set
Mean AUROC of 93.1%, demonstrating strong discrimination
Performance metrics for individual challenge classifiers are detailed in Table A.2. Notably, the system achieved the highest performance on structured audio patterns (playback challenges) and distinctive vocal characteristics (whispering, cup mouth), while performance was comparatively lower on challenges involving subtle modulations in amplitude and speaking rate.
This architecture significantly outperformed traditional approaches based on hand-crafted features such as MFCC, LFCC, or Mel spectrogram representations, validating our choice of the Wav2Vec2 framework as the backbone for challenge compliance detection.
| Challenge | LCNN | SpecR-Net | Star-GANv2-VC | PPG-VC | NISQA-Human | Human-Human |
@@ A.4. Challenge Compliance Detection Architecture
Pitch Comparison, Generalization and Human Agreement. Table A.3 evaluates Pitch’s detection performance across multiple approaches and voice conversion systems. The first two columns show AUC scores for traditional deepfake detectors using LFCC features with LCNN (Todisco et al., 2019) classifier (75.1% average) and SpecRNet (Kawa et al., 2022) (86.0% average). The middle columns demonstrate that our challenge-based approach maintains high detection performance against different voice converters: StarGANv2-VC (Li et al., 2021) (88.2% AUC) and PPG-VC (Liu et al., 2021) (88.3% AUC), comparable to our primary FREE-VC (Li et al., 2023b) evaluation. The rightmost columns present correlation analysis between automated quality assessment (NISQA) and human ratings, revealing weak correlation ( average), which indicates that machines and humans assess different speech quality aspects. This finding supports our human-AI collaborative design, as the two evaluation modalities provide complementary information. The strong human-human agreement ( average using Kendall’s coefficient) validates the consistency and reliability of human evaluators in our studies.